3. Solve the following LPP by the two-phase simplex method. (10) 

 Max 1 2 3 Z = x + x − x

 Subject to 4x1 + x2 + x3 = 4

3 2 6 x1 + x2 − x4 =

 x1

, x2

, x3 ≥ 0



1
Expert's answer
2021-03-31T12:07:59-0400

Solution.

May be the condition of the task next:

"z=x_1+x_2-x_3 - max, \newline
4x_1+x_2+x_3=4,\newline
3x_1+2x_2-x_3=6, \newline
x_1,x_2,x_3>=0."

Introduce new variables "x_4, x_5:"

"4x_1+x_2+x_3+x_4=4, \newline
3x_1+2x_2-x_3+x_5=6,"

then "z = -Mx_4-Mx_5=\newline =-M(4-4x_1-x_2-x_3)-M(6-3x_1-2x_2+x_3)=\newline =7Mx_1+Mx_2-10M - max"

Introduce new variable x_0:

"x_0=7x_1+3x_2."

We will have:

"x_0=-10+7x_1+3x_2,"

"x_4=4-4x_1-x_2-x_3,\newline
x_5=6-3x_1-2x_2+x_3."

This plan is not optimal because there are positive elements in the expression for "x_0" . Select a new variable "x_1," then

"x_1=1-\frac{1}{4}x_2-\frac{1}{4}x_3-\frac{1}{4}x_4."

From here

"x_0=-3+\frac{5}{4}x_2-\frac{7}{4}x_3-\frac{7}{4}x_4,"

"x_5=3-\frac{5}{4}x_2+\frac{7}{4}x_3+\frac{3}{4}x_4."

The expression for "x_0" has positive elements, the plan is not optimal. The largest coefficient is near "x_2" , then choose "x_2" as a new variable.

"x_2=\frac{12}{5}+\frac{7}{5}x_3+\frac{3}{5}x_4-\frac{4}{5}x_5,
\newline
x_0=0-x_4-x_5,
\newline
x_1=\frac{2}{5}-\frac{3}{5}x_3-\frac{2}{5}x_4+\frac{1}{5}x_5."

The expression for "x_0"

has not positive elements, the plan is optimal.

We will deduce variables "x_4, x_5" from basis:

"x_0=\frac{14}{5}-\frac{1}{5}x_3,"

"x_1=\frac{2}{5}-\frac{3}{5}x_3,"

"x_2=\frac{12}{5}+\frac{7}{5}x_3."

Optimal plan:

"x_1=\frac{2}{5}, x_2=2\frac{2}{5}, x_3=0."

So, max "z(x)=2\frac{2}{5}."

Answer. max "z(x)=2\frac{2}{5}."